# Efficient Iterative Proximal Variational Inference Motion Planning

### v3

@@ Abstract
Motion planning under uncertainty can be cast as a stochastic optimal control problem where the optimal posterior distribution has an explicit form. To approximate this posterior, this work frames an optimization problem in the space of Gaussian distributions by solving a Variational Inference (VI) in the path distribution space. For linear-Gaussian stochastic dynamics, we propose a proximal algorithm to solve for an optimal Gaussian proposal iteratively. The computational bottleneck is evaluating the gradients with respect to the proposal over a dense trajectory. We exploit the sparse motion planning factor graph and Gaussian Belief Propagation (GBP), allowing for parallel computing of these gradients on Graphics Processing Units (GPUs). We term the novel paradigm as the Parallel Gaussian Variational Inference Motion Planning (P-GVIMP). Building on the efficient algorithm for linear Gaussian systems, we then propose an iterative paradigm based on Statistical Linear Regression (SLR) techniques to solve motion planning for nonlinear stochastic systems, where the P-GVIMP serves as a sub-routine for the linearized time-varying system. We validate the proposed framework on various robotic systems, demonstrating significant speed acceleration achieved by leveraging parallel computation and successful planning solutions for nonlinear systems under uncertainty. An open-sourced implementation is presented at https://github.com/hzyu17/VIMP.
@@ 1 Introduction
Motion Planning is one core decision-making component in autonomous robotic systems [1, 2]. Given an environment, a start configuration and a goal configuration, a motion planner computes a trajectory connecting the two configurations. The trajectory optimization paradigm [3, 4] formulates the motion planning problem as an optimization over all admissible trajectories with dynamics and environment constraints. An ‘optimal’ plan can be obtained by solving the optimization program by minimizing certain optimality indexes, such as time consumption and control energy.
@@
Uncertainties such as sensing [5], actuation noises, and external disturbances [6] arise and affect the quality of the motion plans in the execution phase. To reduce their impacts, motion planning under uncertainties takes into account the uncertainties in the planning phase in their formulations [7]. Stochastic optimal control and probabilistic robotics [8, 9] provide a principled framework to address this problem, where noise is explicitly modeled in robot dynamics, and a statistical optimality index is minimized over the trajectory distribution space.
@@
Gaussian Process Motion Planning (GPMP) paradigm [10, 11] used the linear SDE dynamics model to formulate motion planning problems as probability inference, where the optimal solution is formulated as a posterior probability. The objective in GPMP is solving a trajectory that maximizes this posterior.
@@
The solution obtained from GPMP is still deterministic. Variational Inference (VI) [12, 13] formulates the inference problem into an optimization by minimizing the Kullback–Leibler (KL) divergence between a proposal distribution and the target posterior. VI has been applied to solve for a trajectory distribution in the planning-as-inference formulation [14, 15, 16, 17]. Gaussian Variational Inference Motion Planning (GVIMP) [14, 15] showed satisfying performances in challenging tasks such as planning through narrow gaps [18]. The GVIMP formulation is related to entropy-maximized motion planning and reinforcement learning [19, 20, 21, 22].
@@
Introducing distributional variables increases the dimension of the problem. Fortunately, in motion planning problems, the underlying probabilistic graph enjoys a sparsity pattern [23, 11] which is leveraged in algorithm designs to factorize the target posterior and reduce the time complexity [24, 15]. In this work, we study the proximal point GVI algorithm [25, 26] over the motion planning factor graph, where we show the updates in explicit form and that these gradient computations can be distributed onto the factor levels over the factor graph.
@@
In terms of computation efficiency, similar to deterministic trajectory optimization paradigms [10, 11, 4], the collision-checking step of the algorithm on a dense trajectory is computationally demanding in GVIMP and represents the computational bottleneck [15]. In this work, we further leverage the sparse factor graph to distribute the collision-checking computations in parallel on a GPU. We also deployed the Gaussian Belief Propagation (GBP) [27, 28] algorithm for efficient marginal covariance computation. We term the algorithm the Parallel Gaussian Variational Inference Motion Planning (PGVIMP).
@@
Most existing planning-as-inference paradigms [11, 29, 14] assume linear SDE as the underlying dynamics. Different from these works, we solve the problem for nonlinear stochastic differential equations (SDEs) [30], where we make the standard Gaussian transition assumption widely used in robotics applications [31, 32]. We then perform the Statistic Linear Regression (SLR) iteratively to obtain a linear time-varying (LTV) system in each iteration and perform PGVIMP on the obtained LTV system. This process integrates the sequential Bayesian principles into the GVIMP framework for motion planning tasks.
@@
(a) Proximal Variational Inference. The proximal point algorithm and its stochastic version were first proposed in [25], where the KL-divergence was used as the proximal operator. The connection between the proximal-point and natural gradient descent algorithms was pointed out in [25, 13]. [26] generalized the proximal point variational inference algorithms by replacing the proximal term with general divergence functions. Recently, [33] proposed a proximal gradient algorithm by decomposing the KL-divergence objective into a potential and an entropy term. The above algorithms were applied directly to the joint posterior over all the variables without considering the factor graph structure of the posterior.
(b) Iterative approximations for nonlinear stochastic processes. The techniques based on iterative linearization to approximate nonlinear stochastic processes originated in the nonlinear state estimation literature. Related works are the Extended Kalman Filter (EKF), which leverages the Taylor approximation [34], statistical approximations based on moment matching [34], Laplace approximation [35], and cubature approximation [36]. In [37] the authors first introduced the use of Gauss-Hermite quadrature methods in the nonlinear estimation problems. To the best of the author’s knowledge, this work is the first to solve nonlinear planning problems by integrating iterative approximations and variational inference.
(c) Parallel Motion Planning. Parallelism and GPU acceleration have been applied to motion planners. The works [38, 39, 40, 41] proposed GPU-accelerated sampling-based motion planning algorithms. [42, 43] used GPU for fast collision-checking, which is a bottleneck in motion planning. The authors of [44] applied GPU acceleration in model predictive control algorithm for online control of high-DOF robotic systems. All these works dealt with collision-checking for deterministic states, while in our work, we solve collision-checking in a probabilistic sense, i.e., we compute an expected collision cost with respect to a proposal distribution over a dense trajectory.
In this work, we leverage this sparse factor and the parallel computing on GPUs to accelerate the computation of two key components in the GVI motion planning paradigm: the parallel computation of the collision likelihood and the computation of marginal covariances over the sparse factor graph.
@@
Key Contributions. The contributions of this work are: (1) We proposed a KL-proximal variational inference algorithm for robot motion planning, where we show that the proximal gradient computations in the paradigm can be parallelized by leveraging a splitting structure and the underlying sparse factor graph; (2) A parallel computation implementation on GPU of the KL-proximal algorithms is proposed, where we leveraged parallel computation of the collision costs and Gaussian Belief Propagation (GBP) to compute marginal covariances efficiently; (3) We propose a sequential statistical linearization scheme for motion planning of nonlinear dynamical systems, where we use the parallel KL-proximal algorithm for linear systems as a submodule at each outer iteration.
Planning results are demonstrated on various robot models and comparison studies have been conducted to demonstrate the efficiency improvement.
@@
This work builds upon the author’s prior work [14, 15], where we addressed the issue of computation efficiency, explored the KL-proximal GVI algorithm in the path distribution space for motion planning, and the applicability of the GVIMP approach to nonlinear dynamics.
@@
This paper is organized as follows. Section 2 introduces the background knowledge. Section 3 introduces the KL-proximal GVI algorithm and its structure for motion planning problems. Section 4 introduces the sparse motion planning factor graph and marginal gradient updates for the KL-proximal algorithm. Section 5 introduces the iterative PGVIMP algorithm for nonlinear systems. The proposed framework is illustrated in Section 7 through numerical experiments, followed by a conclusion in Section 8.
@@ 2.1 Planning under Uncertainties as Stochastic Control
We solve the motion planning under uncertainties using the stochastic control framework. For the nonlinear process
@@
[…]
@@
Gaussian assumptions. The state distribution of system (1) follows the Fokker-Planck (FPK) equations, leading to non-Gaussian distributions for the states under the nonlinearity (1). It is hard to solve the FPK for general nonlinearities , and Gaussian assumptions are one popular approximation assumption in the robotics applications, especially in the filtering and estimation [32, 31].
2.2 Gaussian Variational Inference Motion Planning (GVIMP)
Problem (2) admits an equivalent path distribution control formulation that will be introduced in this section.
(a) Control-inference duality. Denote the measure induced by the controlled linear process (7) as ${\mathbb{Q}}$, and the measure induced by the prior process (3) as ${\mathbb{P}}$. By Girsanov’s theorem [45], the following objective is equivalent to (2) [15, 46]
@@ 2.2 Gaussian Variational Inference Motion Planning (GVIMP)
[…]
@@
(b) Linear-Gaussian System Case. When the stochastic process (1) reduces to linear time-varying stochastic system
@@
Consider the above KL-minimization problem (4) with the time discretization
$\mathbf{t}\triangleq\left[t_{0},\dots,t_{N}\right],\;t_{0}=0,\;t_{N}=T.$ (9)
For linear Gaussian prior dynamics (8), the discrete-time path distribution $d{\mathbb{P}}$ over $\mathbf{t}$ is a Gaussian distribution $\mathbf{X}\sim\mathcal{N}({\boldsymbol{\mu}}|\mathbf{K})$. The discrete-time cost factor $e^{-\mathcal{V}}$ is defined as
[…]
@@
${\mathbb{Q}}^{\star}(\mathbf{X})\propto e^{-\frac{1}{2}\lVert\mathbf{X}-{\boldsymbol{\mu}}\rVert_{\mathbf{K}^{-1}}^{2}-\sum_{i=0}^{N}V(X_{i})}\triangleq\tilde{q}^{\star}(\mathbf{X}).$ (10)
@@
Define the negative log probability for the posterior as $\psi(X)\triangleq-\log\tilde{q}^{\star}(X)$. The Variational Inference objective is equivalent to
$\mathcal{J}(q_{\theta})={\rm KL}\left(q_{\theta}(\mathbf{X})\parallel\tilde{q}^{\star}(\mathbf{X})\right)=\mathbb{E}[\psi(X)]-\mathcal{H}(q_{\theta}),$
where $\mathcal{H}(q_{\theta})$ denotes the entropy of the distribution $q_{\theta}$, which promotes the robustness of the motion plan solution. We add an additional temperature parameter, $\hat{T}$, to trade off this robustness, leading to the objective of temperature [14]
$\mathcal{J}(q_{\theta})={\rm KL}\left(q_{\theta}(\mathbf{X})\parallel\tilde{q}^{\star}(\mathbf{X})\right)=\mathbb{E}[\psi(X)]-\hat{T}\mathcal{H}(q_{\theta})$ (12)
3 KL-Proximal GVI Algorithm For Linear-Gaussian Systems
In this section, we derive a proximal point Gaussian Variational Inference algorithm for the GVIMP problem for linear-Gaussian system formulation (11).
@@ 3.1 Proximal Point Gaussian Variational Inference
(a) KL-proximal point Variational Inference. The proximal point, or proximal minimization algorithm in the Euclidean space, is defined by solving the following optimization problem at each step
@@
where $\mathcal{J}$ is the original objective in a minimization problem, $\lVert x-x_{k}\rVert_{2}^{2}$ is a quadratic regularizer centered around the previous iterate $x_{k}$, and $\beta_{k}$ is the step size. The proximal point algorithm is shown to be more stable than vanilla gradient descent, as it can be interpreted as the backward Euler integration of the gradient flow [47, 48].
In the variational inference problems, the original objective is to maximize the ELBO, $\mathcal{L}$. The KL divergence is a better metric candidate than the Euclidean $2$-norm for the regularizer. Replacing the $2$-norm regularizer in (13) by the KL-divergence, we obtain at the KL-proximal point algorithm, which solves the following optimization at each step [49, 25]
@@
(b) Splitting in Variational Motion Planning. To solve the proximal point iteration for general nonlinear ELBOs turned out to be challenging [26]. A commonly-encountered situation in variational inference is that the ELBO can be splitted into $2$ parts, $-\mathcal{L}=\psi_{e}+\psi_{d}$, where the gradients of $\psi_{e}\triangleq\mathbb{E}_{q_{\theta}}\left[\log\left(f_{e}(x)\right)\right]$ are easy to compute, sometimes are directly known in closed-form, e.g., for Gaussian priors and other conjugate models; and the computation of gradients of $\psi_{d}\triangleq\mathbb{E}_{q_{\theta}}\left[\log\left(f_{d}(x)\right)\right]$ is not direct. In order to obtain the updates efficiently, first-order approximations can be performed on $\psi_{d}$, resulting in proximal gradient algorithms [47]. This decomposition is also known as ‘splitting’.
Inspired by this, we derive the splitting structure for the variational inference in the path distribution space for motion planning problems for the first time. We decompose the motion planning ELBO into two different parts
$\begin{split}\!\!\!\mathcal{L}(x,\theta)&=\mathbb{E}_{q_{\theta}}\left[\log\tilde{{\mathbb{Q}}}^{*}\right]-\mathbb{E}_{q_{\theta}}\left[\log q_{\theta}\right]\\ &=\!-\mathbb{E}_{q_{\theta}}\left[\log q_{\theta}+\frac{1}{2}\lVert\mathbf{x}-{\boldsymbol{\mu}}\rVert_{\mathbf{K}^{-1}}^{2}\right]\!-\!\mathbb{E}_{q_{\theta}}\left[V(\mathbf{x})\right],\end{split}$ (15)
@@
$\psi_{e}(x,\theta)$ $\triangleq\mathbb{E}_{q_{\theta}}\left[\log q_{\theta}+\frac{1}{2}\lVert\mathbf{x}-{\boldsymbol{\mu}}\rVert_{\mathbf{K}^{-1}}^{2}\right],\;$ (16a)
$\psi_{d}(x,\theta)$ $\triangleq\mathbb{E}_{q_{\theta}}\left[V(\mathbf{x})\right].$ (16b)
The function $\psi_{e}$ is the negative likelihood of the prior cost-related term and the entropy, which is a conjugate model. The term $\psi_{d}$ is the cost related to collision avoidance, which is non-conjugate. Consider the Gaussian distribution $q_{e}\sim\mathcal{N}({\boldsymbol{\mu}},\mathbf{K})$, then $\psi_{e}$ can be written as the KL-divergence between $q_{\theta}$ and $q_{e}$ up to a constant as
$\psi_{e}(x,\theta)\propto\mathbb{E}_{q_{\theta}}\left(\log q_{\theta}-\log q_{e}\right)={\rm KL}\left(q_{\theta}\parallel q_{e}\right),$
which indicates that the gradients of $\psi_{e}$ with respect to $(\mu_{\theta},\Sigma_{\theta}^{-1})$ have closed-form. For the non-conjugate model, $\psi_{d}$, we approximate this term $\psi_{d}$ to the first order. The splitting and first-order approximation lead to the following KL proximal gradient update at each iteration.
(c) The KL-proximal gradients with splitting. After splitting, we derive the update step of the KL-proximal algorithm. Substituting the terms $\psi_{e}$ and $\psi_{d}$ in (16) into (14), we have
[…]
Taking the first-order gradients with respect to $(\mu_{\theta},\Sigma_{\theta}^{-1})$, we obtain the KL-proximal updates for them. The update rule is summarized in the following Theorem 1, whose proof can be found in A.
@@
We remark that this update rule is accurate in the sense that we directly approximate the expected term $\psi_{d}(\theta)=\mathbb{E}_{q(\mathbf{x}\mid\theta)}\left[\|\mathbf{h}(\mathbf{x})\|_{\Sigma_{\mathrm{obs}}}^{2}\right]$ to the first-order, instead of approximating the integrand inside the expectation.
(d) Connection to Natural Gradient Descent (NGD). The NGD [24, 15] updates can be integrated into the proximal point algorithms with a symmetric KL-divergence being the proximal regularizer [13]. Specifically, the natural gradient descent update step with unit step size is
@@
With a step size $\beta_{k}$, the NGD update is equivalent to changing the proximal operator in (17) from the KL divergence to the symmetric KL divergence.

### v4

@@ Abstract
We cast motion planning under uncertainty as a stochastic optimal control problem, where the optimal posterior distribution has an explicit form. To approximate this posterior, this work frames an optimization problem in the space of Gaussian distributions by solving a Variational Inference (VI) in the path distribution space. For linear-Gaussian stochastic dynamics, a proximal algorithm is proposed to solve for an optimal Gaussian proposal iteratively. The computational bottleneck is evaluating the gradients with respect to the proposal over a dense trajectory. To tackle this issue, the sparse planning factor graph and Gaussian Belief Propagation (GBP) are exploited, allowing for parallel computation of these gradients on Graphics Processing Units (GPUs). We term the novel paradigm the Parallel Gaussian Variational Inference Motion Planning (P-GVIMP). Building on the efficient algorithm for linear Gaussian systems, we then propose an iterative paradigm based on Statistical Linear Regression (SLR) techniques to solve planning problems for nonlinear stochastic systems, where the P-GVIMP serves as a sub-routine for the linearized time-varying system at each iteration. The proposed framework is validated on various robotic systems, demonstrating significant speed acceleration achieved by leveraging parallel computation and successful planning solutions for nonlinear systems under uncertainty. An open-sourced implementation is presented at https://github.com/hzyu17/VIMP.
@@ 1 Introduction
Motion Planning is one core decision-making component in autonomous robotic systems [1, 2]. Given an environment, a start configuration and a goal configuration, a motion planner computes a trajectory connecting the two configurations. The trajectory optimization paradigm [3, 4] formulates the motion planning problem as an optimization over all admissible trajectories with dynamics and environment constraints. An ‘optimal’ plan is obtained by solving an optimization program that minimizes specific optimality criteria, such as time consumption and control energy.
@@
Uncertainties such as sensing [5], actuation noises, and external disturbances [6] arise and affect the quality of the motion plans in the execution phase. To reduce their impacts, motion planning under uncertainties takes into account the uncertainties in the planning phase in their formulations [7]. Stochastic optimal control and probabilistic robotics [8, 9, 10] provide a principled framework to address this problem, where noise is explicitly modeled in robot dynamics, and a statistical optimality index is minimized over the trajectory distribution space.
@@
Gaussian Process Motion Planning (GPMP) paradigm [11, 12] used the linear SDE dynamics model to formulate motion planning problems as probability inference, where the optimal solution is formulated as a posterior probability. The objective in GPMP is to solve a trajectory that maximizes this posterior.
@@
The solution obtained from GPMP is still deterministic. Variational Inference (VI) [13, 14] formulates the inference problem into an optimization by minimizing the Kullback–Leibler (KL) divergence between a proposal distribution and the target posterior. VI has been applied to solve for a trajectory distribution in the planning-as-inference formulation [15, 16, 17, 18]. Gaussian Variational Inference Motion Planning (GVIMP) [15, 16] showed satisfying performance in challenging tasks such as planning through narrow gaps [19]. The GVIMP formulation is related to entropy-maximized motion planning and reinforcement learning [20, 21, 22, 23].
@@
Introducing distributional variables increases the dimension of the problem. Fortunately, in motion planning problems, the underlying probabilistic graph enjoys a sparsity pattern [24, 12] which is leveraged in algorithm designs to factorize the target posterior and reduce the time complexity [16, 25]. In this paper, we propose the proximal point GVI algorithm [26, 27] over the motion-planning factor graph. The iterative update rule is presented in an explicit form, and its computations are naturally distributed across the factor levels of the graph.
@@
Similar to deterministic trajectory optimization paradigms [4, 12], the collision-checking step of the algorithm on a dense trajectory is computationally demanding in GVIMP and represents the computational bottleneck [16]. The sparse factor graph is leveraged in this paper to distribute collision-checking computations in parallel on the GPU. We also deploy the Gaussian Belief Propagation (GBP) algorithm [28] for efficient marginal covariance computation.
@@
Most existing planning-as-inference paradigms [12, 15, 29, 30] assume linear stochastic differential equations (SDEs) as the underlying dynamics. Unlike these works, this study addresses nonlinear SDEs [31], under the standard Gaussian transition assumption widely adopted in robotics applications [32, 33]. We iteratively apply Statistical Linear Regression (SLR) to obtain a linear time-varying (LTV) system at each time, and perform P-GVIMP on the resulting LTV system. The proposed algorithm integrates the sequential Bayesian principles [34] into the GVIMP framework for motion planning tasks.
@@
(a) Proximal Variational Inference. The proximal point algorithm was first proposed in [26], where the KL-divergence was used as the proximal operator for distributions. The connection between the proximal-point and natural gradient descent algorithms was pointed out in [14]. [27] generalized the proximal point variational inference by replacing the proximal term with general divergence functions. Recently, [35] proposed a proximal gradient algorithm by decomposing the KL-divergence objective into a potential and an entropy term. The above algorithms were applied directly to the posteriors without considering their factor graph structure.
(b) Approximations for nonlinear stochastic processes. The iterative linearization techniques for nonlinear systems approximate a nonlinear model with a linear one at each time step. The Extended Kalman Filter (EKF) [36] uses a single Taylor series expansion. Statistical linearization techniques for nonlinear systems in filtering, such as the Unscented Kalman Filter (UKF) [37], offer a powerful alternative to analytical methods like the EKF. These methods, often referred to as sigma-point filters, deterministically sample points from the state’s probability distribution and then propagate these points through the nonlinear function. The mean and covariance of the transformed points are then used to create a linearized model. This approach doesn’t require computing derivatives and can more accurately capture the statistics of the transformed distribution, especially for highly nonlinear systems. To the best of the author’s knowledge, our work is the first to use the sigma-point paradigm to solve nonlinear planning problems.
(c) Parallel Motion Planning. Parallel motion planning has seen concrete GPU and multi-core implementations that significantly accelerate planning. [38] and [39] demonstrated GPU-accelerated collision checking and PRM construction, achieving orders of magnitude speedups. [40] and [41] exploited batch operations to parallelize sampling-based planning. In trajectory optimization, STOMP [7] explicitly evaluated multiple rollouts in parallel. For multi-agent systems, ORCA [42] and parallelized CBS variants [43] enabled scalable multi-robot coordination. Finally, GPU-RRT-Connect [44] extended parallelism to the full RRT pipeline, showing consistent speedups over CPU baselines. All previous studies addressed parallel collision checking for deterministic motion planning, whereas this paper considers collision checking in a probabilistic setting by computing the expected collision cost with respect to a proposal distribution over a dense trajectory.
@@
Key Contributions. The contributions of this research are: (1) We propose a KL-proximal variational inference algorithm for robot motion planning and show that its proximal gradient computations parallelize naturally by exploiting the splitting structure and the underlying sparse factor graph; (2) A parallel KL-proximal algorithms is proposed via computation implementation on GPUs, where parallel collision cost computation and Gaussian Belief Propagation are used to compute marginal factors efficiently; (3) We propose a sequential statistical linear regression scheme for motion planning of nonlinear dynamical systems, where the parallel KL-proximal algorithm is applied to the linearized systems as a sub-module at each outer iteration.
@@
Planning results are demonstrated on various robot models, and comparison studies have been conducted to demonstrate the efficiency improvement. This paper builds upon the authors’ prior studies [15, 16] by addressing computational efficiency, exploring the KL-proximal GVI algorithm in the path-distribution space for motion planning, and demonstrating the applicability of the GVIMP approach to nonlinear dynamics.
@@
This paper is organized as follows. Section 2 introduces the background knowledge. Section 3 introduces the KL-proximal GVI algorithm and its structure for planning problems. Section 4 introduces the sparse factor graph and marginal gradient updates for the KL-proximal algorithm. Section 5 introduces the iterative P-GVIMP algorithm for nonlinear systems. The proposed framework is illustrated in Section 7 through numerical experiments, followed by a conclusion in Section 8.
@@ 2.1 Planning under Uncertainties as Stochastic Control
We address motion planning under uncertainty within a stochastic control framework. For the nonlinear process
@@
where the running cost consists of a state-related cost $\int_{0}^{T}V(X_{t})dt$, a control energy cost $\int_{0}^{T}\frac{1}{2}\|u_{t}\|^{2}dt$, and $\Psi_{T}(x_{T})$ is a state terminal cost. The state costs regulate the desired behaviors, such as collision avoidance. The prior process associated with the process (1) is defined by letting $u_{t}\equiv 0$, leading to
@@
Gaussian assumptions. The state distribution of system (1) follows the Fokker-Planck (FPK) equations, leading to non-Gaussian distributions for the states under the nonlinearity (1). It is hard to solve the FPK for general nonlinearities [31], and Gaussian assumptions are one popular approximation assumption in robotics applications [33, 32].
2.2 Gaussian VI Motion Planning (GVIMP)
This section covers the problem (2) as a path-distribution control problem, and proposes a Gaussian VI solution to it.
Control-inference duality
Denote the measure induced by the controlled linear process (7) as $\mathbb{Q}$}, and the measure induced by the prior process (3) as $\mathbb{P}$}. By Girsanov’s theorem, the following objective is equivalent to (2) [16, 45]
@@ 2.2 Gaussian VI Motion Planning (GVIMP)
where the cost-related functional $\mathcal{V}\triangleq\int_{0}^{T}V(X_{t})dt+\Psi_{T}(X_{T})$, and the measure $\mathbb{Q}$$^{\star}$ is defined as
@@
Linear-Gaussian System Case
When the stochastic process (1) reduces to linear time-varying stochastic system
@@
Consider the above KL-minimization problem (4) with the time discretization $\mathbf{t}\triangleq\left[t_{0},\dots,t_{N}\right],\;t_{0}=0,\;t_{N}=T$. For linear Gaussian prior dynamics (8), the discrete-time path distribution $d{\mathbb{P}}$ over $\mathbf{t}$ is a Gaussian distribution $\mathbf{x}\sim\mathcal{N}({\boldsymbol{\mu}}|\mathbf{K})$. The discrete-time cost factor $e^{-\mathcal{V}}$ is defined as
$e^{-\mathcal{V}}=e^{-\int_{0}^{T}V(X_{t})dt+\Psi_{T}}\approx e^{-\sum_{i=0}^{N-1}V(X_{i})\times\Delta t+\Psi_{T}},$
@@
${\mathbb{Q}}^{\star}(\mathbf{x})\propto e^{-\frac{1}{2}\lVert\mathbf{x}-{\boldsymbol{\mu}}\rVert_{\mathbf{K}^{-1}}^{2}-\sum_{i=0}^{N}V(X_{i})}\triangleq\tilde{q}^{\star}(\mathbf{x}).$
@@
Define the negative log probability for the posterior as $\psi(\mathbf{x})\triangleq-\log\tilde{q}^{\star}(\mathbf{x})$, and the VI objective is rewritten as
$\mathcal{J}(q_{\theta})={\rm KL}\left(q_{\theta}(\mathbf{x})\parallel\tilde{q}^{\star}(\mathbf{x})\right)=\mathbb{E}[\psi(\mathbf{x})]-\mathcal{H}(q_{\theta}),$
where $\mathcal{H}(q_{\theta})$ denotes the entropy of the distribution $q_{\theta}$, which promotes the robustness of the motion plan solution. An additional temperature parameter, $\hat{T}$, is added to trade off this robustness, leading to the objective of temperature [15]
$\mathcal{J}(q_{\theta})={\rm KL}\left(q_{\theta}(\mathbf{x})\parallel\tilde{q}^{\star}(\mathbf{x})\right)=\mathbb{E}[\psi(\mathbf{x})]-\hat{T}\mathcal{H}(q_{\theta})$
3 KL-Proximal Algorithm For Linear-Gaussian Systems
This section derives a proximal point algorithm for the GVIMP problem for linear systems (9).
@@ 3.1 Proximal Point Gaussian Variational Inference
(a) KL-proximal point Variational Inference. At each step, the proximal point, or proximal minimization algorithm in the Euclidean space, solves the following optimization problem
@@
where $\lVert x-x_{k}\rVert_{2}^{2}$ is a regularizer centered around the previous iterate $x_{k}$, and $\beta_{k}$ is the step size. The proximal point algorithm is more stable than vanilla gradient descent, since it corresponds to the backward Euler integration of the gradient flow [46].
The KL divergence is a better metric candidate regularizer than the Euclidean $2$-norm for distributions. Replacing the $2$-norm regularizer in (10) by the KL-divergence, the KL-proximal point algorithm [47] is obtained at each step
@@
(b) Splitting in Variational Motion Planning. Solving the proximal point iteration for general nonlinear Evidence Lower Bounds (ELBOs) is challenging [27]. A common situation in variational inference is that the ELBO splits into two parts, $\mathcal{L}=\psi_{e}+\psi_{d}$, where the gradients of $\psi_{e}\triangleq\mathbb{E}_{q_{\theta}}\left[\log\left(f_{e}(x)\right)\right]$ are easy to compute, sometimes are directly known in closed form, e.g., for Gaussian priors and other conjugate models; and the computation of gradients of $\psi_{d}\triangleq\mathbb{E}_{q_{\theta}}\left[\log\left(f_{d}(x)\right)\right]$ is not direct. To obtain updates efficiently, the first-order approximations are performed on $\psi_{d}$. This decomposition is also known as ‘splitting’.
Inspired by this, this paper derives the splitting structure for motion planning problems for the first time. We decompose the motion planning ELBO into two different parts
$\begin{split}\!\!\!\mathcal{L}(\mathbf{x},\theta)&=\mathbb{E}_{q_{\theta}}\left[\log\tilde{{\mathbb{Q}}}^{*}\right]-\mathbb{E}_{q_{\theta}}\left[\log q_{\theta}\right]\\ &=\!-\mathbb{E}_{q_{\theta}}\left[\log q_{\theta}+\frac{1}{2}\lVert\mathbf{x}-{\boldsymbol{\mu}}\rVert_{\mathbf{K}^{-1}}^{2}\right]\!-\!\mathbb{E}_{q_{\theta}}\left[V(\mathbf{x})\right],\end{split}$ (12)
@@
$\psi_{e}(\mathbf{x},\theta)$ $\triangleq\mathbb{E}_{q_{\theta}}\left[\log q_{\theta}+\frac{1}{2}\lVert\mathbf{x}-{\boldsymbol{\mu}}\rVert_{\mathbf{K}^{-1}}^{2}\right],\;$ (13a)
$\psi_{d}(\mathbf{x},\theta)$ $\triangleq\mathbb{E}_{q_{\theta}}\left[V(\mathbf{x})\right].$ (13b)
The function $\psi_{e}$ is the negative likelihood of the prior cost-related term and the entropy, which is a conjugate model. The term $\psi_{d}$ is the cost related to collision avoidance, which is non-conjugate. Let $q_{e}\sim\mathcal{N}({\boldsymbol{\mu}},\mathbf{K})$. Then $\psi_{e}$ equals the KL divergence between $q_{\theta}$ and $q_{e}$, up to a constant
$\psi_{e}(\mathbf{x},\theta)\propto\mathbb{E}_{q_{\theta}}\left(\log q_{\theta}-\log q_{e}\right)={\rm KL}\left(q_{\theta}\parallel q_{e}\right),$
which indicates that the gradients of $\psi_{e}$ with respect to $(\mu_{\theta},\Sigma_{\theta}^{-1})$ have closed-form. The non-conjugate model, $\psi_{d}$, is approximated to the first order. This leads to the following KL proximal gradient update at each iteration.
(c) The KL-proximal gradients with splitting. After splitting, the update step of the KL-proximal algorithm is derived. By substituting the terms $\psi_{e}$ and $\psi_{d}$ from (13) into (11), we obtain
$\theta_{k+1}=\underset{\theta}{\operatorname{argmin}}\;\mathcal{J}_{\rm ProxKL}$ (14)
$\triangleq\underset{\theta}{\operatorname{argmin}}\;\theta^{T}\left[\nabla\psi_{d}\left(\theta_{k}\right)\right]+\psi_{e}(\theta)+\frac{{\rm KL}\left(q_{\theta}(\mathbf{x})\|q\left(\mathbf{x}\mid\theta_{k}\right)\right)}{\beta_{k}}.$
Taking the first-order gradients with respect to $(\mu_{\theta},\Sigma_{\theta}^{-1})$, we obtain the KL-proximal updates for them. The update rule is summarized in Theorem 1. See A for the proof.
@@
This update rule is accurate in the sense that it directly approximates the expected term $\psi_{d}(\theta)=\mathbb{E}_{q(\mathbf{x}\mid\theta)}\!\left[\|\mathbf{h}(\mathbf{x})\|^{2}_{\Sigma_{\mathrm{obs}}}\right]$ to first order, rather than approximating the integrand inside the expectation. Under $L$-smooth assumption on $\psi_{d}$, the approximation error is of order $\mathcal{O}(\|\theta-\theta_{k}\|^{2})$ and hence remains bounded. B presents the proof.
Convergence of the algorithm. First assume that the ELBO $\mathcal{L}$ is continuous and admits a maximum. Based on [27], the following assumptions are required: (i) $\psi_{d}$ is $L$-smooth; (ii) $\psi_{e}$ is convex; (iii) There exists $\alpha>0$ satisfying the specific monotonicity condition. Under these assumptions, the proposed KL-Proximal update satisfies $\|\theta_{k+1}-\theta_{k}\|\to 0$ when $0<\beta_{k}<\alpha/L$. B provides a detailed statement and proof.
(d) Connection to Natural Gradient Descent (NGD). The NGD updates can be integrated into proximal point algorithms, using the symmetric KL-divergence as the proximal regularizer [14]. Specifically, the NGD update step with unit step size is
@@
With step size $\beta_{k}$, the NGD update replaces the KL proximal term in (14) with its symmetrized counterpart.